The first 1st,5th and 10th term of a linear sequence are geometric progression. If the 2nd and 8th term of the linear sequence is 30, find the non-zero common difference of the linear sequence



1
Expert's answer
2021-09-13T13:36:02-0400

The 2nd and 8th term of the linear sequence is 30

"a+d+a+7d = 30 \\

2a + 8d = 30 \\

a + 4d = 15 \\

a = 15-4d"

1st, 5th and 10th terms of a linear sequence are geometric progression

"\frac{(a+4d)}{a} = \frac{(a+9d)}{(a+4d)} \\

(a+4d)^2 = a(a+9d) \\

a^2 + 8ad + 16d^2 = a^2 + 9ad \\

16d^2 = ad \\

16d^2 = d(15-4d) \\

16d^2 = 15d – 4d^2 \\

20d^2 - 15d = 0 \\

5d(4d - 3) = 0"

d = 0 or "d = \frac{3}{4}"

if d = 0, then a = 15, and the linear sequence would be

15, 15, 15

if "d= \frac{3}{4}" , "a = 15 - 4(\frac{3}{4}) = 12"

the sequence would be

"12, \frac{51}{4}, \frac{27}{2}, ..."

sum of 2nd and 8th

"= \frac{51}{4} + 12+7(\frac{3}{4}) = \frac{51}{4} + \frac{69}{4} = 30"

term1 = 12

term5 "= 12+4(\frac{3}{4}) = 15"

term10 "= 12 + 9(\frac{3}{4}) = \frac{75}{4}"

term5/term1 "= \frac{15}{12} = \frac{5}{4}"

term10/term5 "= \frac{(\frac{75}{4}) }{(15)} = \frac{5}{4}"